What is the magnitude of the electric force, F in
Answer #1

Q1 = +25 nC

Q2 = -36 nC

d = 35 cm

a) F = kQ1Q2/d^2 = (9*10^9*25*10^-9*36*10^-9)/0.35^2 = 6.61*10^-5 N

b) Direction of F (attractive) is towards Q1 (-ve x-direction)

c) Acceleration = a = F/m = 66.1*10^-6/8.5*10^-9 = 7.78*10^3 m/s